# Cryptanalysis on Lightweight Verifiable Homomorphic Encryption

### v2

@@ 4.3 Attack on Replication Encoding with Multiple Secret Keys
First, let the adversary fix a random subset $A\subset\mathbb{Z}_{t}$ where the message $m$ is expected to belong. Next, evaluate $\chi_{A}$ homomorphically on a ciphertext. After evaluating $\chi_{A}$, it becomes possible to forge the values that exactly belong to $A$. If $\{m,v_{i}:i\in S\}\cap A=\{m\}$, then ${\sf CMult}$ between ${}_{S}:=\textsf{Eval}_{\chi_{A}}(\textsf{ct}{})$ ct ) and the encoding of the vector $(\overbrace{1,\dots,1}^{n},\overbrace{0,\dots,0}^{N-n})$ is an encryption of a vector $(J_{S^{c}}\,\|\,\overbrace{0\dots 0}^{N-n})$.
@@
4:     ${}_{\textsf{bool}}\leftarrow\textsf{Eval}_{\chi_{A}}(\textsf{ct}{})$ ct )
@@
Let us calculate the attack success probability with respect to the size of $A$. Since $A$ is chosen randomly, the success probability of the attack depends only on the size of $A$. Let $p=|A|/{t}$ be the probability of each element $m$ or $v_{i}$ belonging to the set $A$. Set $q:=1-p$ and $\mu:=\left|\{m,v_{i}:i\in S\}\right|=1+\frac{n}{2}$.
@@
Let $A$ be a randomly chosen subset of $\mathbb{Z}_{t}$ with size $|A|=p\cdot t$. For $\textsf{ct}{}_{S^{c}}\leftarrow\mathsf{Extended\_Attack}(\textsf{ct}{},A)$ in the Algorithm 4, the probability
@@
Pseudo-Random Characteristic Function on BFV.
The attack cost in Theorem 4.3 is determined by the evaluation cost of the characteristic function. In the attack on vADDG, the evaluation was straightforward since extracting encrypted bits from the TFHE ciphertext string was easy. However, in BFV, homomorphically evaluating $\chi_{A}$ for any randomly given subset $A$ with the full support $\mathbb{Z}_{t}$ generally requires $O(\sqrt{t})$ homomorphic operations, which results in an exponential cost with respect to $\log t$. Therefore, it is necessary to devise an efficient method for evaluating a pseudorandom characteristic function over $\mathbb{Z}_{t}$.
Before constructing a pseudo-random characteristic function, we note that the pseudorandom characteristic function used in this attack does not need to be cryptographically secure; rather, it is sufficient for it to be a function that heuristically produces an unbiased uniform distribution.
To construct an efficient pseudo-random characteristic function, we assume the pseudo-randomness of the distribution of the roots of unity in $\mathbb{Z}_{t}$: For an integer $t$ and $d\mid\phi(t)$, let $U_{t,d}:=\{x\in\mathbb{Z}_{t}^{\times}:x^{\phi(t)/d}=1\}\subset\mathbb{Z}_{t}^{\times}$. We assume that this subset $U_{t,d}$ of size $\phi(t)/d$ is an unbiased sample from a uniform distribution of $(\phi(t)/d)$-combinations from $\mathbb{Z}_{t}^{\times}$. Specifically, for a large odd prime $p$, $U_{p,2}$ is the set of quadratic residues modulo $p$, whose pseudo-randomness is utilized to construct the Legendre PRF [Dam90].
Now we state the property of the characteristic function $\chi_{U_{t,d}}:\mathbb{Z}_{t}\to\{0,1\}$ for $t=p^{r}$. Before proof, note that $a\in\mathbb{Z}_{p^{r}}$ is a zero divisor if and only if $a\equiv 0\pmod{p}$.
@@
    First, suppose that $x$ is a non-unit. Then the left-hand side $(\sf LHS)$ is zero. Also as $x$ is a multiple of $p$, the right-hand side $(\sf RHS)$ is also 00 since it would be multiple of $p^{\phi(t)}\equiv 0\pmod{p^{r}}$.
@@
    Next, suppose that $x\in U_{t,d}$. Since $x$ is a unit, $x^{\phi(t)}=1$. Also $x^{\phi(t)/d}=1$ by definition. Thus both $\sf LHS$ and $\sf RHS$ are equal to $d$.
@@
    Lastly, suppose that $x\in\mathbb{Z}_{t}^{\times}\setminus U_{t,d}$. Since $x$ is a unit, $x^{\phi(t)}=1$. Also $x^{\phi(t)/d}\neq 1$ by definition, while $(x^{\phi(t)/d})^{d}-1=(x^{\phi(t)/d}-1)\left(\sum_{i=0}^{d-1}\left(x^{\phi(t)/d}\right)^{i}\right)=0$. Thus, $\sum_{i=0}^{d-1}\left(x^{\phi(t)/d}\right)^{i}=0$ or both $x^{\phi(t)/d}-1,\sum_{i=0}^{d-1}\left(x^{\phi(t)/d}\right)^{i}$ are zero-divisors. We exclude the second case to conclude that $\sf RHS=0$: For $y=x^{\phi(t)/d}$, if $y-1$ is a zero divisor, then $y=pm+1$ for some integer $m$, consequently $\sum_{i=0}^{d-1}\left(x^{\phi(t)/d}\right)^{i}\equiv d\not\equiv 0\pmod{p}.$ Thus it is not a zero divisor, a contradiction.
Note that if $t$ is a prime, then the $x^{\phi(t)}$ term in RHS is not needed. Also, note that this function can be homomorphically evaluated using $O(\log t)$ homomorphic multiplications with a multiplication depth of $\lceil\log\phi(t)\rceil$.
By homomorphically evaluating $\chi_{U_{t,d}}(x+a)$ for a random integer $a$ and appropriate $d\mid(p-1)$ with asymptotic size $d\approx\mu$, the adversary can construct a characteristic function $\chi_{A}$ with an asymptotic size of $|A|=\phi(t)/d\approx\lfloor t/\mu\rceil$ to achieve the maximal attack probability in the Theorem 4.3. If necessary, the adversary may evaluate several characteristic functions to construct another size of characteristic function. [^3] For example, one may utilize the formula $\chi_{A}(x)\cdot\chi_{B}(x)=\chi_{A\cap B}(x).$
@@ 5.1 Polynomial Encoding
There is another suggested scheme for verifiable HE in [CKP+24], the Polynomial Encoding (PE). Its plaintext space and ciphertext space are BFV plaintext/ciphertext space but with a new indeterminate $Y$. More precisely, its encryption is
@@
where the $\textsf{Enc}_{\textsf{PE}}{}$ is the coefficient-wise BFV encryption. To avoid confusion, we represent the elements of $\mathcal{R}^{2}_{q}[Y]$ using capitalized sans-serif font: $\mathsf{F}(Y)=\textsf{ct}{}_{0}+\textsf{ct}{}_{1}Y+\dots+\textsf{ct}{}_{d}Y^{d}$. We refer to $\mathsf{F}(Y)$ as a ciphertext polynomial and each of its coefficients as a ciphertext.
Encryption in PE proceeds as follows: For the number of slots $N$, a random verification value $\mathbf{v}=(v_{1},\dots,v_{N})\in\mathbb{Z}_{t}^{N}$ and a secret value $\alpha\in\mathbb{Z}_{t}^{\times}$ are chosen. Then, the message $\mathbf{m}=(m_{1},\dots,m_{N})\in\mathbb{Z}_{t}^{N}$ and the verification value $\mathbf{v}$ are interpolated at $Y=0$ and $Y=\alpha$, respectively. After that, each message is encrypted coefficient-wisely:
[…]
The server coefficient-wisely performs homomorphic operations like addition, constant multiplication, rotation, and even bootstrapping, except for multiplication. The multiplication is evaluated as a polynomial in $Y$: For example, the multiplication between two ciphertext polynomials $\textsf{ct}{}_{0}+\textsf{ct}{}_{1}Y$ and $\textsf{ct}{}_{2}+\textsf{ct}{}_{3}Y$ is
@@ 5.2 Attack on Polynomial Encoding
4:     $\textsf{ct}{}_{\sf forged}\leftarrow\textsf{Eval}_{g}(\mathsf{L}_{1}(0),\dots,\mathsf{L}_{k}(0))$triangleright $g$ is any circuit different from $f$.
5:     $\mathsf{L}^{\star}(Y)\leftarrow\textsf{Enc}(u_{1},\dots,u_{N})Y$triangleright Every $u_{i}$ is a unit.
6:     $\mathsf{P}(Y)\leftarrow p({\sf L}^{\star}(Y))$triangleright $p(y)=y^{\phi(t)}$.
@@ 6.2 Attack Results
The experiments were performed on an Intel(R) Xeon(R) Silver 4114 CPU at 2.20GHz running Linux in a single-threaded environment. The attack results are summarized in Tables 33. The time cost only measured evaluation time for $\textsf{Enc}(J_{S^{c}})$.
@@
According to Theorem 4.2, the expected attack success probabilities given by $1-\binom{32}{k}/\binom{64}{k}$ are $50\%$, $75.40\%$, $94.34\%$, and $100\%$ for $k=1$, $2$, $4$, and $33$, respectively. This attack circuit can be implemented without bootstrapping under the parameters $(\log N=15,\log PQ=880,\log Q=759,t=2^{16}+1)$. As shown in Table 3, the attack time increases linearly with $k$. We note that Algorithm 3 is inherently parallelizable, as it performs ${\sf CVS}(\textsf{ct}{},i)$ and ${\sf is\_in\_set}(\textsf{ct}{},i)$ independently for each $i=1,\dots,k$. Thus, by parallelizing the execution of ${\sf CVS}(\textsf{ct}{},i)$, the overall attack time can be reduced to under one minute even with $k=33$. In addition, leveraging GPU-based optimizations could potentially reduce the runtime to just a few seconds. This assumption is realistic, given that adversaries in the context of HE are typically assumed to have access to significant computational resources.
@@
Since $\phi(t)=2^{16}$, when constructing the characteristic function in Lemma 2, we can choose $d$ to be a power of 2. In Experiment 5, where $n=64$, we set $d=32\cong\mu=33$ for the characteristic function. As a result, the output becomes a scalar multiple of $\textsf{Enc}(J_{S^{c}})$ with non-negligible probability. According to Theorem 4.3, the corresponding expected attack success probability is $\left(1-\frac{2^{11}}{2^{16}+1}\right)^{32}-\left(1-\frac{2^{11}}{2^{16}+1}\right)^{33}\cong 1.13\%$. Also in Experiment 6, where $n=32$, we set $d=16\cong\mu=17$. The corresponding expected attack success probability is $\left(1-\frac{2^{12}}{2^{16}+1}\right)^{16}-\left(1-\frac{2^{12}}{2^{16}+1}\right)^{17}\cong 2.22\%$.
Parameters $\log N$ $\log PQ$ $\log Q$ $t$ $|S|=|\{1,\dots,n\}|$ $k$
REP Attack Exp1 15 880 759 $12^{16}+$ 1 64 1
@@
Evaluation Time (Sec/op) Attack Time (Evaluating $\textsf{Enc}(J_{S^{c}})$) Remark
REP Attack Exp1 8.37s $\textsf{rtk}{}_{1}$ generated.
@@
Probability Theoretical Attack Success Probability Experimental Attack Success Probability # of Iterations
REP Attack Exp1 50.00% 51.8% 1000
@@
Due to the lack of BFV bootstrapping functionality in the Lattigo library, we implemented the attacks on REP and $\textsf{REP}{}^{\sf msk}$ only for specific parameter settings, and did not implement the attack on PE. However, we note that in a bootstrappable setting, not only is an attack on PE feasible, but the above attacks can also be carried out independently of the parameter settings.
@@ 7.1 Homomorphic Cryptanalysis
To provide verifiability, the VHE schemes in [CKP+24, ADDG24] utilized auxiliary secret information as well as the secret key of the HE: a set of indices $S$ in REP and an element $\alpha\in\mathbb{Z}_{t}^{\times}$ in PE. Attacks against the schemes described in this paper consist of two steps as follows;
@@
For PE, if $\textsf{Enc}(\alpha)$ is recovered we can modify a legitimate ciphertext easily. However, ReQ adopted a re-randomized process to make recovering $\textsf{Enc}(\alpha)$ hard. Here the proposed attack does not learn the phase of learning $\alpha$. This is not only because, as mentioned earlier, PE encoding can be performed without knowledge of $\alpha$ or $\textsf{Enc}(\alpha)$, but it can also be explained from the following perspective: We can think of the ciphertext polynomial $\mathsf{F}(Y)\in\mathcal{R}^{2}_{q}[Y]$ as a ciphertext with two secret keys
@@
is an encryption of $\alpha$ under $(1,s,\alpha,\alpha s)$ and an encryption of 00 under $(1,s,0,0)$. Indeed, one of the possible choice of $\mathsf{L}^{\star}(Y)$ is $\textsf{Enc}(0,\dots,0)+\textsf{Enc}(1,\dots,1)Y$, which is as an encryption of $\alpha$ and 00 in this very sense.

### v3

@@ 4.3 Attack on Replication Encoding with Multiple Secret Keys
First, let the adversary fix a random subset $A\subset\mathbb{Z}_{t}$ where the message $m$ is expected to belong. Next, evaluate $\chi_{A}$ homomorphically on a ciphertext. After evaluating $\chi_{A}$, it becomes possible to forge the values that exactly belong to $A$. If $\{m,v_{i}:i\in S\}\cap A=\{m\}$, then $\sf CMult$} between $\textsf{ct}{}_{S}:=\textsf{Eval}_{\chi_{A}}(\textsf{ct}{})$ and the encoding of the vector $(\overbrace{1\dots 1}^{n}\|\overbrace{0\dots 0}^{N-n})$ is an encryption of a vector $(J_{S^{c}}\,\|\,\overbrace{0\dots 0}^{N-n})$.
@@
4:  $\textsf{ct}{}_{\textsf{bool}}\leftarrow\textsf{Eval}_{\chi_{A}}(\textsf{ct}{})$
@@
Let us calculate the attack success probability with respect to the size of $A$. Since $A$ is chosen randomly, the success probability of the attack depends only on the size of $A$. Let p=|A|/tp=|A|/{t} be the probability of each element $m$ or $v_{i}$ belonging to the set $A$. Set q:=1pq:=1-p and $\mu:=\left|\{m,v_{i}:i\in S\}\right|=1+\frac{n}{2}$.
@@
Let $A$ be a randomly chosen subset of $\mathbb{Z}_{t}$ with size $|A|=p\cdot t$. For $\textsf{ct}{}_{S^{c}}\leftarrow\mathsf{Extended\_Attack}(\textsf{ct}{},A)$ in Algorithm 4, the probability
@@
Pseudorandom Characteristic Function over $\mathbb{Z}_{t}$.
The attack cost in Theorem 4.3 is determined by the evaluation cost of the characteristic function. In the attack on vADDG, the evaluation was straightforward since extracting encrypted bits from the TFHE ciphertext string was easy. However, homomorphically evaluating $\chi_{A}$ for any randomly given subset $A$ with the full support $\mathbb{Z}_{t}$ generally requires $O(\sqrt{t})$ homomorphic operations, which results in an exponential cost with respect to $\log t$. Therefore, it is necessary to devise an efficient (polynomial time) method for evaluating a pseudorandom characteristic function over $\mathbb{Z}_{t}$.
Before constructing a pseudorandom characteristic function, note that the pseudorandom characteristic function used in this attack does not need to be cryptographically secure; rather, it is sufficient for it to be a function that heuristically produces an unbiased uniform distribution.
To construct an efficient pseudorandom characteristic function, we assume the pseudorandomness of the distribution of the roots of unity in $\mathbb{Z}_{t}$: For an integer $t$ and $d\mid\phi(t)$, let $U_{t,d}:=\{x\in\mathbb{Z}_{t}^{\times}:x^{\phi(t)/d}=1\}\subset\mathbb{Z}_{t}^{\times}$. We assume that this subset $U_{t,d}$ of size $\phi(t)/d$ is an unbiased sample from a uniform distribution of $(\phi(t)/d)$-combinations from $\mathbb{Z}_{t}^{\times}$. Specifically, for a large odd prime $p$, $U_{p,2}$ is the set of quadratic residues modulo $p$, whose pseudorandomness is utilized to construct the Legendre PRF [Dam90]. Although the specific assumptions and parameters in here are different from those of Legendre PRF, we again note that it is sufficient for it to be a function that heuristically produces an unbiased uniform sampling.
Now we state the property of the characteristic function $\chi_{U_{t,d}}:\mathbb{Z}_{t}\to\{0,1\}$ for $t=p$. For $t=p^{r},r>1$, refer Appendix 0.B.
@@
    First, suppose that $x=0$. Then the left-hand side $(\sf LHS)$ and the right-hand side $(\sf RHS)$ is 0.
@@
    Next, suppose that $x\in U_{t,d}$. Since $x^{\phi(t)/d}=1$ by definition of $U_{t,d}$, both $\sf LHS$ and $\sf RHS$ are equal to $d$.
@@
    Lastly, suppose that $x\in\mathbb{Z}_{t}^{\times}\setminus U_{t,d}$. Then, $x^{\phi(t)/d}\neq 1$ by definition of $U_{t,d}$ Also, $(x^{\phi(t)/d})^{d}-1=(x^{\phi(t)/d}-1)\left(\sum_{i=0}^{d-1}\left(x^{\phi(t)/d}\right)^{i}\right)=0$. Thus, $\sum_{i=0}^{d-1}\left(x^{\phi(t)/d}\right)^{i}=0$ since $\mathbb{Z}_{t}$ is an integral domain.
Note that this function can be homomorphically evaluated using $O(\log t)$ homomorphic multiplications with a multiplication depth of $\lceil\log\phi(t)\rceil$.
By homomorphically evaluating $\chi_{U_{t,d}}(x+a)$ for a random integer $a$ and an appropriate divisor $d\mid(p-1)$ with $d\approx\mu$, the adversary can construct a characteristic function $\chi_{A}$ whose support has size $|A|=\phi(t)/d\approx\lfloor t/\mu\rceil$, thereby achieving the maximal attack probability in Theorem 4.3. If necessary, the adversary may evaluate several characteristic functions to construct another size of characteristic function. [^4] For example, one may utilize the formula $\chi_{A}(x)\cdot\chi_{B}(x)=\chi_{A\cap B}(x).$
@@ 5.1 Polynomial Encoding
There is another suggested scheme for verifiable HE in [CKP+24], the Polynomial Encoding (PE). Its plaintext and ciphertext spaces are inherited from the original homomorphic encryption scheme, augmented with a new indeterminate $Y$. For example, in the BFV scheme, its encryption is
@@
where the $\textsf{Enc}_{\textsf{PE}}{}$ is the coefficient-wise encryption. To avoid confusion, we represent the elements of $\mathcal{R}^{2}_{q}[Y]$ using capitalized sans-serif font: $\mathsf{F}(Y)=\textsf{ct}{}_{0}+\textsf{ct}{}_{1}Y+\dots+\textsf{ct}{}_{d}Y^{d}$. We refer to $\mathsf{F}(Y)$ as a ciphertext polynomial and each of its coefficients as a ciphertext.
Encryption in PE proceeds as follows: For the number of slots $N$, a random verification value $\mathbf{v}=(v_{1},\dots,v_{N})\in\mathbb{Z}_{t}^{N}$ and a secret value $\alpha\in\mathbb{Z}_{t}^{\times}$ are chosen. Then, the message $\mathbf{m}=(m_{1},\dots,m_{N})\in\mathbb{Z}_{t}^{N}$ and the verification value $\mathbf{v}$ are interpolated at $Y=0$ and $Y=\alpha$, respectively. After that, each message is encrypted coefficientwise:
$\mathbb{Z}^{N}_{t}$$\mathbb{Z}^{N}_{t}[Y]$$\mathcal{R}^{2}_{q}[Y]$$\mathbf{m}$$\mathbf{m}+\left[\frac{\mathbf{v-m}}{\alpha}\right]_{t}Y$$\textsf{ct}{}_{0}+\textsf{ct}{}_{1}Y$$\textsf{Enc}_{\textsf{PE}}{}$$\textsf{Enc}_{\textsf{PE}}{}$
The server coefficientwise performs homomorphic operations like addition, constant multiplication, rotation, and even bootstrapping, except for multiplication. The multiplication is evaluated as a polynomial in $Y$: For example, the multiplication between two ciphertext polynomials $\textsf{ct}{}_{0}+\textsf{ct}{}_{1}Y$ and $\textsf{ct}{}_{2}+\textsf{ct}{}_{3}Y$ is
@@ 5.2 Attack on Polynomial Encoding
4:  $\textsf{ct}{}_{\sf forged}\leftarrow\textsf{Eval}_{g}(\mathsf{L}_{1}(0),\dots,\mathsf{L}_{k}(0))$ $\triangleright$ $g$ is any circuit different from $f$.
5:  $\mathsf{L}^{\star}(Y)\leftarrow\textsf{Enc}(u_{1},\dots,u_{N})Y$ $\triangleright$ Every $u_{i}$ is a unit.
6:  $\mathsf{P}(Y)\leftarrow p({\sf L}^{\star}(Y))$ $\triangleright$ $p(y)=y^{\phi(t)}$.
@@ 6.2 Attack Results
The experiments were performed on an Intel(R) Xeon(R) Silver 4114 CPU at 2.20GHz running Linux in a single-threaded environment. The attack results are summarized in Tables 3, 3. The time cost only measured evaluation time for $\textsf{Enc}(J_{S^{c}})$.
@@
According to Theorem 4.2, the expected attack success probabilities given by $1-\binom{32}{k}/\binom{64}{k}$ are $50\%$, $75.40\%$, $94.34\%$, and $100\%$ for $k=1$, 22, 44, and $33$, respectively. This attack circuit can be implemented without bootstrapping under the parameters $(\log N=15,\log PQ=880,\log Q=759,t=2^{16}+1)$. As shown in Table 3, the attack time increases linearly with $k$, but we note that Algorithm 3 is inherently parallelizable, as it performs ${\sf CVS}(\textsf{ct}{},i)$ and ${\sf is\_in\_set}(\textsf{ct}{},i)$ independently for each $i=1,\dots,k$. Thus, by parallelizing the execution of ${\sf CVS}(\textsf{ct}{},i)$, the overall attack time can be reduced to under one minute even with $k=33$. In addition, leveraging GPU-based optimizations could potentially reduce the runtime to just a few seconds. This assumption is realistic, given that adversaries in the context of HE are typically assumed to have access to significant computational resources.
@@
Since $\phi(t)=2^{16}$, when constructing the characteristic function in Lemma 1, we can choose $d$ to be a power of 2. In Experiment 5, where $n=64$, we set $d=32\cong\mu=33$ for the characteristic function. As a result, the output becomes a scalar multiple of $\textsf{Enc}(J_{S^{c}})$ with non-negligible probability. According to Theorem 4.3, the corresponding expected attack success probability is $\left(1-\frac{2^{11}}{2^{16}+1}\right)^{32}-\left(1-\frac{2^{11}}{2^{16}+1}\right)^{33}\cong 1.13\%$. Also in Experiment 6, where $n=32$, we set $d=16\cong\mu=17$. The corresponding expected attack success probability is $\left(1-\frac{2^{12}}{2^{16}+1}\right)^{16}-\left(1-\frac{2^{12}}{2^{16}+1}\right)^{17}\cong 2.22\%$. The parameters and results are presented in Table 3,3 and 3.
Parameters $\log N$ $\log PQ$ $\log Q$
$t$
$|S|=|\{1,\dots,n\}|$
$k$
REP Attack Exp1 15 880 759 $2^{16}+1$ 64 1
@@
Evaluation Time
(Sec/op)
Attack Time
(Evaluating $\textsf{Enc}(J_{S^{c}})$)
Remark
REP Attack
Exp1 8.37s
$\textsf{rtk}{}_{1}$ generated.
@@
Probability
Theoretical Attack
Success Probability
Experimental Attack
Success Probability
# of Iterations
REP Attack
Exp1 50.00% 51.8% 1000
@@
We remark that the drop in success probability observed in Exp. 5 and Exp. 6 is within one standard deviation. Also, due to the lack of BFV bootstrapping functionality in the Lattigo library, we implemented the attacks on REP and $\textsf{REP}{}^{\sf msk}$ only for specific parameter settings, and did not implement the attack on PE. However in a bootstrappable setting, not only is an attack on PE feasible, but the above attacks could also be carried out independently of the parameter settings.
@@ 7.1 Homomorphic Cryptanalysis
To provide verifiability, the schemes considered in this paper utilize auxiliary secret information in addition to the secret key of the HE scheme: a set of indices $S$ in REP (which also applies essentially to vADDG) and an element $\alpha\in\mathbb{Z}_{t}^{\times}$ in PE. Attacks against the schemes described in this paper consist of two steps as follows;
@@
For PE, if $\textsf{Enc}(\alpha)$ is recovered we can modify a legitimate ciphertext easily. However, ReQ adopted a re-randomized process to make recovering $\textsf{Enc}(\alpha)$ hard. Here the proposed attack does not learn $\alpha$. This is not only because, as mentioned earlier, PE encoding can be performed without knowledge of $\alpha$ or $\textsf{Enc}(\alpha)$, but it can also be explained from the following perspective: We can think of the ciphertext polynomial $\mathsf{F}(Y)\in\mathcal{R}^{2}_{q}[Y]$ as a ciphertext with two secret keys
@@
is an encryption of $\alpha$ under $(1,s,\alpha,\alpha s)$ and an encryption of 0 under $(1,s,0,0)$. Indeed, one of the possible choice of $\mathsf{L}^{\star}(Y)$ is $\textsf{Enc}(0,\dots,0)+\textsf{Enc}(1,\dots,1)Y$, which is as an encryption of $\alpha$ and 0 in this very sense.